A rocket is launched with velocity 10km/s. If radius of earth is R, then maximum height attained by it will be
 
(a) 2R
(b)3R
(c)4R
(d)5R
 
Please help.

Grade:11

1 Answers

Sitanshu kumar Singh
6 years ago
If a body is projected with velocity v which is less than escape velocity for earth which is 11.2 km/s. then maximum height attained by the body is you can find it by this formula h=R(radius of earth)/[v`(escape velocity )/v(velocity with which you project)2-1].If you didn`t understand the formula let me know I will explain it. h=R/(11.2/10)2-1=R/ 0.2544=4R(approx)